The perimeter of a rectangle is 60 cm. If the ratio of two adjacent sides is 3:2, find the length and width of the rectangle.
Answer by TimothyLamb(4379)

x = length
y = width
---
2x + 2y = 60
x/y = 3/2
2x = 3y
---
put the system of linear equations into standard form
---
2x + 2y = 60
2x - 3y = 0
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 18 cm
y = 12 cm
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---